Calculate the mass of water that can be produced when 85 g CH4 reacts with 155 g O2 according to the equation: CH4(g) + 2O2(g) → CO2(g) + 2H2O(g)
Answer #1

moto mon (My = 12 +4=16900 1602 = 12 +32 = 44 g surf O2 = 32g/mol f120= 18 9hnol Bare DTID CH4 (9) + 20249) = (029) + 24209)thank you very much.hope this will help you.good luck